### v1

# Molecular Seeds of Shear: An operator-level necessity result for first-order Chapman–Enskog deviatoric stress

Abstract
We provide an explicit functional-analytic formulation and rigorous derivation showing that, in closed and unforced kinetic systems under Chapman–Enskog scaling, anisotropic (deviatoric) stress appears only when the first Chapman–Enskog kinetic correction $f^{(1)}$ is nonzero. Under precise nullspace, coercivity, and Fredholm solvability assumptions on the linearized collision operator, and with uniform remainder control, we prove that a vanishing first correction implies the absence of $O(\varepsilon)$ deviatoric stress in the hydrodynamic limit. A fully worked BGK example and remainder estimates are included. We discuss implications for how microscopic seeds (deterministic departures or finite-$N$ fluctuations) project into macroscopic amplification channels for transition and turbulence. The result is proved under explicit nullspace, coercivity (or hypocoercivity) and Fredholm solvability hypotheses for the linearized collision operator, together with uniform remainder bounds; these assumptions are stated precisely in Section 4 and Appendix A and restrict the class of admissible interaction kernels to those listed in A1.
1 Introduction
Turbulence is a macroscopic manifestation of nonlinear instability and multiscale energy transfer, yet its molecular antecedents remain conceptually unsettled when compared to continuum operator-level theory. The Boltzmann equation for the single-particle distribution $f(x,v,t)$ furnishes the kinetic bridge between molecular dynamics and continuum fluid models. The Chapman–Enskog expansion systematically expresses macroscopic fluxes and stresses as consequences of small deviations of $f$ from a local Maxwellian equilibrium [1, 2].
This manuscript strengthens that classical connection by providing explicit functional-analytic hypotheses and a rigorous derivation that all anisotropic (deviatoric) stress appears at first order in the Chapman–Enskog expansion. Under precise nullspace, coercivity, and Fredholm solvability assumptions on the linearized collision operator, and in the absence of external forcing or inhomogeneous walls, we prove that an initial Chapman–Enskog correction $f^{(1)}\equiv 0$ yields vanishing $O(\varepsilon)$ deviatoric stress in the hydrodynamic limit; consequently, leading-order shear production and its energetic pathways are absent unless a nonzero kinetic correction is present. This deterministic result complements hydrodynamic-limit programs by isolating the operator bounds and solvability conditions required to obtain viscous constitutive relations from kinetic theory [3, 4, 5].
We also quantify how extremely small kinetic departures from equilibrium—whether produced by finite-$N$ statistical fluctuations, weak external perturbations, or minute deterministic inhomogeneities—project onto macroscopic perturbation energy and thereby enter linear and nonmodal amplification channels well studied in hydrodynamic stability theory [6, 7]. Embedding these amplification mechanisms within a Chapman–Enskog framework equipped with explicit operator bounds and remainder estimates supplies a testable pathway from microscopic kinetic deviations to macroscopic instability and transition.
2 Previous work
The Chapman–Enskog procedure, first systematically presented in [1] and summarized in modern monographs [2], provides the classical asymptotic route from the Boltzmann equation to the compressible and incompressible Navier–Stokes equations. Rigorous mathematical formulations of hydrodynamic limits and related compactness arguments were developed by Bardos, Golse, and Levermore [3, 4] and organized in Saint-Raymond’s monograph [5].
Simplified kinetic models such as the BGK model are widely used for explicit calculations of transport coefficients and constitutive laws [8, 9]. Fluctuating hydrodynamics and stochastic kinetic analyses (Landau–Lifshitz formalism and subsequent kinetic-level stochastic equations) quantify how molecular noise projects onto hydrodynamic modes [10, 11]. Independent advances in fluid-mechanics stability theory (nonmodal transient growth) demonstrate that nonnormal linearized operators can produce large short-time energy amplification from tiny seeds [7, 6].
Despite this progress, relatively few studies have made the functional-analytic properties of the Boltzmann collision operator the explicit linchpin connecting kinetic solvability conditions to the onset of macroscopic shear and turbulent energetics. This paper addresses that gap by specifying operator-level hypotheses (nullspace, coercivity, Fredholm solvability), deriving the mapping from the Chapman–Enskog first correction $f^{(1)}$ to the deviatoric stress, and explaining how microscopic seeds are necessary kinetic precursors for $O(\varepsilon)$ shear production in closed, unforced systems.
3 Model and notation
We consider the Boltzmann equation for a dilute monoatomic gas in $d$ dimensions,
$\partial_{t}f+v\cdot\nabla_{x}f=Q(f,f),\qquad(x,v)\in\Omega_{x}\times\mathbb{R}^{d},\;t>0,$ (3.1)
where $f(x,v,t)$ is the single-particle molecular distribution function and $Q(f,f)$ denotes the Boltzmann collision operator conserving mass, momentum, and energy [1, 2]. For relevant interaction laws the operator may be linearized about a local Maxwellian $M[\rho,u,T]$, yielding the linearized collision operator
Lg:=DQ[M]g,Lg:=DQ[M]\,g, (3.2)
which acts on weighted velocity spaces such as $L^{2}_{v}(M^{-1}\,dv)$ or polynomially weighted $L^{2}_{v}(\langle v\rangle^{k}dv)$ and admits the standard nullspace of collision invariants under mild hypotheses [3, 4, 5].
Macroscopic fields are velocity moments of $f$,
$\rho(x,t)=\int_{\mathbb{R}^{d}}f\,dv,\qquad\rho u(x,t)=\int_{\mathbb{R}^{d}}vf\,dv,\qquad E(x,t)=\int_{\mathbb{R}^{d}}\tfrac{1}{2}|v|^{2}f\,dv,$ (3.3)
and the pressure tensor is the centered second moment,
$P_{ij}(x,t)=\int_{\mathbb{R}^{d}}(v_{i}-u_{i})(v_{j}-u_{j})\,f\,dv,$ (3.4)
with deviatoric stress
$\tau_{ij}=P_{ij}-\tfrac{1}{d}P_{kk}\,\delta_{ij}.$ (3.5)
For analytic transparency we use the BGK relaxation model as an instructive example:
$Q_{\mathrm{BGK}}(f)=\frac{1}{\tau}\bigl(M[f]-f\bigr),$ (3.6)
with relaxation time $\tau$ and local Maxwellian $M[f]$ determined by the moments of $f$ [8, 9]. The Chapman–Enskog expansion is taken under the small Knudsen scaling $\varepsilon=\mathrm{Kn}\ll 1$:
$f=f^{(0)}+\varepsilon f^{(1)}+\varepsilon^{2}f^{(2)}+O(\varepsilon^{3}).$ (3.7)
At leading order $f^{(0)}=M[\rho,u,T]$ and satisfies the compressible Euler equations; the first correction $f^{(1)}$ supplies the viscous stress and heat flux whose constitutive forms reproduce Navier–Stokes transport coefficients after inversion of the appropriate linearized operator [1, 2, 8].
Throughout we adopt the operator-level assumptions used in the core arguments in Section 4 and the notation $N(L)$ for the nullspace of $L$, $L^{-1}$ for the pseudoinverse on $N(L)^{\perp}$, and $\langle\cdot,\cdot\rangle_{L^{2}_{v}(M^{-1})}$ for the velocity-space inner product.
4 Assumptions (A1–A8) — functional-analytic hypotheses
A1 (Collision kernel and mapping properties).
The collision operator $Q(f,f)$ corresponds to a physically realistic molecular interaction law (hard spheres or cutoff Maxwell-type potentials) for which $Q$ is well defined as a bilinear map
$Q:L^{1}_{v}(\langle v\rangle^{m})\times L^{1}_{v}(\langle v\rangle^{m})\to L^{1}_{v}(\langle v\rangle^{m})$ (4.1)
and, under the Maxwellian weight, as a bounded bilinear form on velocity-weighted $L^{2}$ spaces. The operator satisfies conservation identities and the H-theorem [1, 2].
A2 (Spatial domain and boundary conditions).
Either (i) $\Omega_{x}=\mathbb{T}^{d}$ (periodic box), or (ii) $\Omega_{x}=\mathbb{R}^{d}$ with sufficient spatial decay so surface terms vanish. When $\Omega_{x}=\mathbb{R}^{d}$ we require sufficient decay of macroscopic fields and kinetic perturbations so that all integrations by parts performed in the energy estimates produce no boundary terms; concretely we assume $(\rho,u,T)$ and the kinetic remainder lie in $H^{s}_{x}(\mathbb{R}^{d})$ with $s>d/2$ and polynomial velocity weights controlling the moments used below. Inhomogeneous walls are excluded in the main deterministic theorem.
A3 (Initial data and moment/regularity class).
Initial data satisfy
$f_{0}-M\in H^{s}_{x}L^{2}_{v}(\langle v\rangle^{k}),\qquad s>d/2,\;k\geq 2,$ (4.2)
for a global Maxwellian $M$ and weights chosen to control required moments.
A4 (Chapman–Enskog scaling and remainder control).
With $\varepsilon$ the Knudsen number we seek expansions (3.7) with $f^{(0)}=M[\rho,u,T]$ and remainder $R_{\varepsilon}$ satisfying a uniform bound
$\|R_{\varepsilon}\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\leq C\varepsilon^{2}$ (4.3)
for $0<\varepsilon<\varepsilon_{0}$ under the stated regularity and coercivity hypotheses. When not provable for a chosen kernel, the remainder estimate is stated explicitly.
A5 (Spectral and coercivity properties of $L$).
Fix a smooth local Maxwellian $M(x,v)=M[\rho(x),u(x),T(x)]$ with macroscopic fields in $H^{s}_{x}$. Let $L=DQ[M]$ act on $L^{2}_{v}(M^{-1}dv)$. Then:
(i) $N(L)=\mathrm{span}\{1,v_{1},\dots,v_{d},|v|^{2}\}$ and $L^{2}_{v}(M^{-1})=N(L)\oplus N(L)^{\perp}$.
(ii) Coercivity/spectral gap: either there exists $\lambda_{0}>0$ such that for all $g\in N(L)^{\perp}$,
$\langle-Lg,g\rangle_{L^{2}_{v}(M^{-1})}\geq\lambda_{0}\|g\|_{L^{2}_{v}(M^{-1})}^{2},$ (4.4)
or, when a uniform gap is unavailable, $L$ satisfies quantitative hypocoercive estimates yielding analogous control with velocity-derivative terms and associated constants [12, 14, 15].
For concreteness we allow hard-sphere and cutoff Maxwell-type kernels. When we assert a uniform spectral gap we mean the coercivity bound (4.4) holding with a constant $\lambda_{0}>0$ uniform in $x$ for these kernels, as established in the references collected in Appendix A. When a uniform gap is unavailable we instead invoke the quantitative hypocoercivity bounds referenced in A5, with the precise derivative-loss and weighted Sobolev spaces used later specified in Appendix A.
A6 (Fredholm solvability and bounded inverse on $N(L)^{\perp}$).
For each $x$, $Lg=h$ admits a solution $g\in N(L)^{\perp}$ iff $h$ is orthogonal to $N(L)$. Concretely, the pseudoinverse $L_{x}^{-1}$ is taken as the bounded inverse on the orthogonal complement $N(L_{x})^{\perp}$ in the weighted space $L^{2}_{v}(M(x,\cdot)^{-1}\,dv)$ (or the corresponding weighted Sobolev space when hypocoercivity is invoked). We assume the uniform operator bound $\|L_{x}^{-1}\|\leq C_{\mathrm{inv}}$ for all $x$ under the macroscopic regularity A7; Appendix A records the constructive dependence of $C_{\mathrm{inv}}$ on spectral-gap or hypocoercivity constants and on Maxwellian parameters. The pseudoinverse $L^{-1}:N(L)^{\perp}\to N(L)^{\perp}$ depends continuously on smooth variations of $M(x,v)$ [14, 16].
A7 (Regularity in $x$ of macroscopic fields).
The fields $\rho,u,T$ that define $M(x,v)$ lie in $H^{s}_{x}$ with $s>d/2$ and possess uniformly bounded derivatives up to the order required by the hypocoercivity estimates. This guarantees $v\cdot\nabla_{x}f^{(0)}$ belongs to the velocity-weighted spaces on which $L^{-1}$ acts.
A8 (Closed, unforced system).
The kinetic evolution is deterministic and closed: no external body forces, volumetric stochastic forcing, or inhomogeneous boundary injections enter the equation in the main theorem. Finite-$N$ stochastic effects and forced or wall-driven shear are discussed separately.
5 Functional-analytic lemmas and solvability
We collect the operator facts used in the Chapman–Enskog inversion and remainder estimates.
Lemma 5.1 (Nullspace and coercivity).
Under A1 and A5, the linearized collision operator $L=DQ[M]$ satisfies $N(L)=\mathrm{span}\{1,v_{1},\dots,v_{d},|v|^{2}\}$ and the coercivity estimate (4.4) (or the hypocoercive alternative) holds pointwise in $x$ for $g\in N(L)^{\perp}$.
Lemma 5.2 (Fredholm solvability and bounded pseudoinverse).
Under A6 and the coercivity/hypocoercivity hypotheses of Lemma 5.1, for each $x$ the equation $Lg=h$ admits a solution $g\in N(L)^{\perp}$ iff $\int h\phi\,dv=0$ for all collision invariants $\phi$. Moreover there exists a bounded pseudoinverse $L^{-1}:N(L)^{\perp}\to N(L)^{\perp}$ with uniform bound $\|L^{-1}\|\leq C_{\mathrm{inv}}$ under the regularity bounds of A7.
As a consequence, whenever the solvability conditions hold pointwise in $x$ we recover the Chapman–Enskog first correction in the form
$f^{(1)}(x,v)=-\,L^{-1}\bigl((\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}\bigr)(x,v),$ (5.1)
with mapping bounded in the prescribed norms.
Lemma 5.3 (Projection identity and admissibility of streaming).
Let $P$ denote the orthogonal projection in $L^{2}_{v}(M^{-1})$ onto $N(L)$. If $f^{(0)}(x,v)=M[\rho(x),u(x),T(x)]$ is an exact local Maxwellian whose macroscopic moments satisfy the leading-order conservation laws (Euler equations) pointwise in $x$, then
$P\bigl((\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}\bigr)=0.$ (5.2)
The identity (5.2) requires that the macroscopic fields $(\rho,u,T)$ defining $f^{(0)}$ satisfy the leading-order Euler conservation laws pointwise. If the Euler equations hold only up to $O(\varepsilon)$ compatibility errors then the projection produces commutator/source terms of size $O(\varepsilon)$; such terms are tracked in the remainder estimates of Section B and absorbed by the Grönwall closure under the smallness threshold $\varepsilon<\varepsilon_{0}$ specified in Section 6. Hence the streaming/source term lies in $N(L)^{\perp}$ and inversion by $L^{-1}$ is admissible pointwise in $x$.
6 Main theorem
We now state the principal result.
Theorem 6.1 (Molecular seed of shear).
Assume A1–A8 and the operator bounds of Lemmas 5.15.3. Let $\{f_{\varepsilon}\}_{0<\varepsilon<\varepsilon_{0}}$ be a family of sufficiently regular solutions of the rescaled Boltzmann equation under Chapman–Enskog scaling on the domain specified in A2, and write the Chapman–Enskog expansion
$f_{\varepsilon}(x,v,t)=f^{(0)}(x,v,t)+\varepsilon f^{(1)}(x,v,t)+R_{\varepsilon}(x,v,t),$ (6.1)
with $f^{(0)}(x,v,t)=M[\rho,u,T]$. Assume the remainder satisfies the uniform bound
$\|R_{\varepsilon}\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\leq C\varepsilon^{2}$ (6.2)
for $0<\varepsilon<\varepsilon_{0}$ under the norms and regularity in A3–A7. All assertions of Theorem 6.1 are local-in-time on any interval $[0,T]$ on which the macroscopic fields remain in $H^{s}_{x}$ with the uniform bounds of A7. Constants appearing in (6.2)–(6.6) depend explicitly on $\|(\rho,u,T)\|_{H^{s}_{x}}$, the spectral-gap or hypocoercivity constants, the uniform pseudoinverse bound $C_{\mathrm{inv}}$, and the domain geometry. Uniqueness of $f^{(1)}$ is to be understood pointwise in $x$ in the chosen weighted space under the orthogonality conditions (6.5); potential initial-layer corrections that violate (6.5) are treated as part of the parallel remainder and are controlled by the estimates in Section B. Then:
  1. (i)
    (Constitutive decomposition) The pressure tensor admits the asymptotic decomposition
    $P(x,t)=p(x,t)\,I+\varepsilon\,\tau^{(1)}(x,t)+\mathcal{O}(\varepsilon^{2})$ (6.3)
    in $L^{2}_{x}$, where
    $\tau^{(1)}_{ij}(x,t)=\int_{\mathbb{R}^{d}}(v_{i}-u_{i})(v_{j}-u_{j})\,f^{(1)}(x,v,t)\,dv.$ (6.4)
    The mapping $f^{(1)}\mapsto\tau^{(1)}$ is linear and bounded under the prescribed velocity-moment weights.
  2. (ii)
    (Vanishing for spatially uniform Maxwellians) If $f^{(0)}(x,v,t)$ is spatially uniform (so that $(\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}\equiv 0$ pointwise in $x$) and the solvability/Fredholm conditions of A6 hold, then $f^{(1)}(x,v,t)\equiv 0$ for all $(x,t)$ and consequently $\tau^{(1)}\equiv 0$. Hence, in this closed, unforced setting there is no $O(\varepsilon)$ deviatoric stress emerging from the Chapman–Enskog expansion.
  3. (iii)
    (Uniqueness and representation) Conversely, whenever the solvability orthogonality
    $\int_{\mathbb{R}^{d}}\bigl((\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}\bigr)\,\phi(v)\,dv=0$ (6.5)
    holds for all collision invariants $\phi(v)$, the first correction is uniquely determined pointwise in $x$ by the bounded pseudoinverse:
    $f^{(1)}(x,v,t)=-\,L_{x}^{-1}\bigl((\partial_{t}^{(0)}+v\cdot\nabla_{x})f^{(0)}\bigr)(x,v,t),$ (6.6)
    where $L_{x}$ denotes the linearized operator at the local Maxwellian $M[\rho(x,t),u(x,t),T(x,t)]$. The map $f^{(0)}\mapsto f^{(1)}$ is continuous in the weighted norms specified in A3–A7 with operator norm controlled by the inversion constant $C_{\mathrm{inv}}$ from Lemma 5.2.
Corollary 6.2 (Deterministic necessity of kinetic correction for $O(\varepsilon)$ shear).
Under the hypotheses of Theorem 6.1, production of $O(\varepsilon)$ deviatoric stress in a closed, unforced kinetic system necessarily requires a nonzero first-order Chapman–Enskog correction $f^{(1)}$.
7 BGK worked example
For the BGK model
$\partial_{t}f+v\cdot\nabla_{x}f=\frac{1}{\tau}(M[f]-f),$ (7.1)
seek $f=M+\tau f^{(1)}+O(\tau^{2})$. The $O(1)$ equation yields $f^{(0)}=M$. At $O(\tau)$ we have
$(\partial_{t}+v\cdot\nabla_{x})M=-f^{(1)}.$ (7.2)
Evaluating the contribution of $f^{(1)}$ to the pressure tensor at a point $x_{0}$, applying a Galilean shift so $u(x_{0})=0$, and using standard Gaussian integrals for a Maxwellian
$M(v)=\frac{\rho}{(2\pi RT)^{d/2}}\exp\!\bigl(-\tfrac{|v|^{2}}{2RT}\bigr),$ (7.3)
we use moments
$\int v_{i}v_{j}M(v)\,dv=\rho RT\,\delta_{ij},$ (7.4)
$\int v_{i}v_{j}v_{k}v_{l}\,M(v)\,dv=\rho(RT)^{2}(\delta_{ij}\delta_{kl}+\delta_{ik}\delta_{jl}+\delta_{il}\delta_{jk}).$ (7.5)
Then
$\tau^{(1)}_{ij}=\int v_{i}v_{j}f^{(1)}\,dv=-\int v_{i}v_{j}(\partial_{t}+v\cdot\nabla_{x})M\,dv,$ (7.6)
and collecting terms yields the viscous constitutive relation
$\tau^{(1)}_{ij}=-2\mu\,S_{ij},\qquad\mu=\rho\,\tau\,R\,T,$ (7.7)
with $S_{ij}=\tfrac{1}{2}(\partial_{i}u_{j}+\partial_{j}u_{i})-\tfrac{1}{d}(\nabla\cdot u)\,\delta_{ij}$. The Gaussian integrals and prefactors are evaluated under the Galilean shift $u(x_{0})=0$. The BGK remainder estimate is valid in the same weighted $L^{2}_{v}(M^{-1})$ space and the constant in the $O(\varepsilon^{2})$ bound depends explicitly on $\tau$ and $\|(\rho,u,T)\|_{H^{s}_{x}}$.
8 Energy budget, linear stability and transient growth
We summarize the connection between kinetic seed amplitude and macroscopic amplification.
8.1 Kinetic-energy balance
Multiplying the macroscopic momentum equation by $u$ and integrating yields (compressible form suppressed)
$\frac{d}{dt}\int\tfrac{1}{2}\rho|u|^{2}\,dx=-\int\tau_{ij}\partial_{i}u_{j}\,dx+W_{\mathrm{ext}}+B_{\mathrm{boundary}}.$ (8.1)
Substituting the constitutive law $\tau=-\mu S(u)$ recovers viscous dissipation.
8.2 Linearization about a shear profile
For incompressible Navier–Stokes with base flow $U(y)\mathbf{e}_{x}$ and perturbations $u^{\prime}$, the linearized equation is
$\partial_{t}u^{\prime}+U(y)\partial_{x}u^{\prime}+v^{\prime}U^{\prime}(y)\mathbf{e}_{x}=-\nabla p^{\prime}+\nu\Delta u^{\prime},$ (8.2)
with perturbation energy equation
$\frac{d}{dt}E(t)=-\int u^{\prime}_{x}u^{\prime}_{y}U^{\prime}(y)\,dx\,dy\,dz-\nu\int|\nabla u^{\prime}|^{2}\,dV,\qquad E(t)=\tfrac{1}{2}\int|u^{\prime}|^{2}\,dV.$ (8.3)
The production term admits transient amplification for suitable phase relations between $u^{\prime}_{x}$ and $u^{\prime}_{y}$.
8.3 Non-normality and transient growth
The linear operator is non-self-adjoint; the operator norm $\|e^{Lt}\|$ can be large for finite $t$ even if spectral abscissa is negative. Define
$G_{\max}=\sup_{t\geq 0}\|e^{Lt}\|_{L^{2}\to L^{2}}^{2},$ (8.4)
and note $G_{\max}$ can scale like $\mathrm{Re}^{2}$ for some flows [6, 7].
8.4 Linking seed amplitude to threshold
Let $a_{0}$ denote amplitude of the relevant seed projection (e.g., $f^{(1)}$ mapped to initial perturbation energy). After amplification the amplitude is approximately $G^{1/2}a_{0}$; transition to nonlinear cascade is plausible when $G^{1/2}a_{0}\gtrsim a_{\mathrm{NL}}$.
9 Remainder estimates and Chapman–Enskog error bounds
We state a model-precise remainder estimate.
Theorem 9.1 (Remainder estimate — model form).
Let A1–A7 hold and let $\{f_{\varepsilon}\}_{0<\varepsilon<\varepsilon_{0}}$ be solutions admitting
$f_{\varepsilon}=f^{(0)}+\varepsilon f^{(1)}+R_{\varepsilon},$ (9.1)
with $f^{(1)}$ given by (5.1) wherever solvability holds. Assume:
  1. (H1)
    Spectral/coercivity hypothesis: either a uniform spectral gap $\lambda_{0}>0$ is available or $L$ satisfies quantitative hypocoercivity estimates [12, 14, 15].
  2. (H2)
    Uniform macroscopic regularity: $\rho,u,T\in H^{s}_{x}$ with $s>d/2$ and uniform bounds controlling $L$ and $L^{-1}$.
  3. (H3)
    Compatibility and moment control: initial data satisfy A3 and Chapman–Enskog solvability compatibility at order $\varepsilon$, and initial remainders are $O(\varepsilon^{2})$.
Under H1–H3 there exist constants
$C=C\big(T,\|(\rho,u,T)\|_{H^{s}_{x}},C_{\mathrm{inv}}\big),\qquad\varepsilon_{0}=\varepsilon_{0}\big(\|(\rho,u,T)\|_{H^{s}_{x}},C_{\mathrm{inv}}\big),$
such that for every $0<\varepsilon<\varepsilon_{0}$ and every $t\in[0,T]$ the remainder satisfies the explicit bound
$\|R_{\varepsilon}(t)\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\leq C\,\varepsilon^{2}.$
All constants appearing in the energy inequalities of Appendix B are traced to the coercivity/hypocoercivity constants and the $H^{s}_{x}$ norms of the macroscopic fields; in particular the threshold $\varepsilon_{0}$ is determined by the ratio of dissipation to commutator/source constants appearing in the estimates.
Sketch of proof.
Subtracting equations for $f^{(0)}$ and $f^{(1)}$ from the kinetic equation yields an evolution for $R_{\varepsilon}$ of the form
$\partial_{t}R_{\varepsilon}+v\cdot\nabla_{x}R_{\varepsilon}-\frac{1}{\varepsilon}LR_{\varepsilon}=\mathcal{S}_{\varepsilon},$ (9.2)
where $\mathcal{S}_{\varepsilon}$ collects higher-order residuals. Projecting onto $N(L)^{\perp}$ and using coercivity/hypocoercivity yields dissipation estimates; moment equations control projections on $N(L)$. Energy estimates with a Grönwall closure (detailed in Appendix B) yield the uniform $O(\varepsilon^{2})$ bound under H1–H3 [1, 2, 3, 4, 14, 15, 16].
10 Conclusion
Under explicit functional-analytic hypotheses we prove a deterministic operator-level mechanism by which first-order Chapman–Enskog kinetic corrections produce viscous constitutive behavior and enable macroscopic deviatoric stress in closed, unforced systems. The theorem asserts that production of $O(\varepsilon)$ deviatoric stress requires a nonzero $f^{(1)}$. A constructive BGK calculation reproduces the classical viscosity coefficient and demonstrates inversion and remainder control in an explicit setting. Extensions to boundary-driven or forced problems require boundary-layer analysis and are outlined for future work.
Acknowledgements
The author gratefully acknowledges Dr. Spyksma for his insightful and thorough review of the manuscript and Dr. Adam Barkman for his careful proofreading.
Appendix A Pseudoinverse construction: full statement and proof sketch
This appendix provides a detailed statement and proof sketch of the pseudoinverse construction used in the main text.
Proposition A.1 (Pseudoinverse construction — full).
Let $M(x,v)=M[\rho(x),u(x),T(x)]$ with $\rho,u,T\in H^{s}_{x}$, $s>d/2$, uniformly bounded and sufficiently smooth. Assume $L_{x}=DQ[M(x,\cdot)]$ satisfies the uniform spectral-gap condition: there exists $\lambda_{0}>0$ such that for all $x$ and all $g\in N(L_{x})^{\perp}$,
$\langle-L_{x}g,g\rangle_{v,M}\geq\lambda_{0}\|g\|_{L^{2}_{v}(M^{-1})}^{2}.$
Then for each $x$ the restriction $L_{x}|_{N(L_{x})^{\perp}}$ is invertible and the pseudoinverse
$L_{x}^{-1}:N(L_{x})^{\perp}\to N(L_{x})^{\perp}$
is bounded with
$\|L_{x}^{-1}h\|_{L^{2}_{v}(M^{-1})}\leq C_{\mathrm{inv}}\|h\|_{L^{2}_{v}(M^{-1})},$
where $C_{\mathrm{inv}}$ depends only on $\lambda_{0}$ and uniform bounds on $M$ (and not on $\varepsilon$). Additionally $x\mapsto L_{x}^{-1}$ is continuous in operator norm provided $\rho,u,T$ vary continuously in $H^{s}_{x}$.
Proof sketch.
We give a step-by-step constructive argument and indicate precise references where the black-box results are invoked.
Step 1: Frozen-$x$ operator and spectral gap. Fix $x_{0}$. Consider the frozen operator $L_{x_{0}}$ acting on $L^{2}_{v}(M(x_{0},\cdot)^{-1}dv)$. By hypothesis the spectral gap $\lambda_{0}>0$ on $N(L_{x_{0}})^{\perp}$ holds; therefore the restriction $L_{x_{0}}:N(L_{x_{0}})^{\perp}\to N(L_{x_{0}})^{\perp}$ is boundedly invertible. This is a standard consequence of coercivity; see Baranger & Mouhot [13, Thm. 1] or Mouhot & Strain [14] for constructive statements in kinetic settings.
Step 2: Uniformity in $x$ via continuous dependence. The Maxwellian $M(x,v)$ depends smoothly on $(\rho,u,T)(x)$. Under $H^{s}_{x}$ control ($s>d/2$) pointwise values of macroscopic fields vary continuously and uniformly in $x$; the mapping $x\mapsto L_{x}$ is continuous in operator norm on the chosen velocity-weighted space. Using perturbation theory for linear operators (Kato [18]), a spectral gap lower bound persists under small operator perturbations and the resolvent depends continuously on the perturbation. Since the gap is uniform by assumption, there exists a uniform inverse bound $C_{\mathrm{inv}}$ valid for all $x$ in the considered compact set (or uniformly if global bounds apply). For explicit constructive dependence of the coercivity constant on the Maxwellian parameters, see Mouhot & Strain [14] and Baranger & Mouhot [13].
Step 3: Construction of $L_{x}^{-1}$ on $N(L_{x})^{\perp}$. For each $x$ define the orthogonal projector $P_{x}$ onto $N(L_{x})$ in $L^{2}_{v}(M(x,\cdot)^{-1}dv)$ and write $L_{x}=L_{x}|_{N(L_{x})^{\perp}}\oplus 0$ on the decomposition. The inverse is given on $N(L_{x})^{\perp}$ by Riesz theory; the coercivity lower bound provides $\|L_{x}^{-1}\|\leq\lambda_{0}^{-1}$ modulo norm equivalence constants depending on weights (again see [14]).
Step 4: Continuity in operator norm. Since $x\mapsto L_{x}$ is continuous and the inversion map is continuous on the set of invertible operators with uniform gap, $x\mapsto L_{x}^{-1}$ is continuous in operator norm. This step uses standard resolvent identity manipulations and Kato-type estimates; see Kato [18].
Remark on hypocoercivity: If the spectral gap hypothesis is replaced by a hypocoercivity framework (A5(b)), one obtains a bounded inverse in hypocoercive norms (with derivative losses) by following Villani’s constructive Lyapunov functional approach [15] and constructive implementations in [14]. The main difference is (i) $L^{-1}$ is bounded on a higher-regularity velocity space $H^{s_{v}}_{v}(M^{-1})$ and (ii) constants reflect derivative gains/losses required to close the hypocoercive estimates. ∎
Appendix B Remainder estimate: detailed proof (Grönwall closure)
This appendix provides the detailed energy estimates for the remainder $R_{\varepsilon}$ used in the main theorem.
B.1 Equation for the remainder
Subtract the Chapman–Enskog truncation from the kinetic equation to obtain the PDE for $R_{\varepsilon}$:
$\partial_{t}R_{\varepsilon}+v\cdot\nabla_{x}R_{\varepsilon}-\frac{1}{\varepsilon}L_{x}R_{\varepsilon}=\mathcal{S}_{\varepsilon},$
where $\mathcal{S}_{\varepsilon}$ collects quadratic/higher-order terms and explicit Taylor remainders in $\varepsilon$ arising from the nonlinear collision operator and time dependence of $f^{(0)}$ and $f^{(1)}$.
B.2 Projection and splitting
Let $P_{x}$ be the orthogonal projector onto $N(L_{x})$ and split
$R_{\varepsilon}=P_{x}R_{\varepsilon}+(I-P_{x})R_{\varepsilon}=:R_{\varepsilon}^{\parallel}+R_{\varepsilon}^{\perp}.$
Projecting the remainder equation onto $N(L_{x})^{\perp}$ yields
$\partial_{t}R_{\varepsilon}^{\perp}+v\cdot\nabla_{x}R_{\varepsilon}^{\perp}-\frac{1}{\varepsilon}L_{x}R_{\varepsilon}^{\perp}=(I-P_{x})\mathcal{S}_{\varepsilon}+\mathcal{C}_{\varepsilon},$
where $\mathcal{C}_{\varepsilon}$ denotes commutator terms arising from $\partial_{t},P_{x}]$ and $v\cdot\nabla_{x},P_{x}]$ that are controlled using macroscopic regularity. In particular, under A7 one shows (for suitable weighted norms)
$\|\mathcal{C}_{\varepsilon}\|\leq C_{\mathrm{comm}}\Big(\|R_{\varepsilon}^{\perp}\|+\|R_{\varepsilon}^{\parallel}\|\Big),$
with $C_{\mathrm{comm}}$ depending on $\|(\rho,u,T)\|_{H^{s}_{x}}$ and derivatives appearing in the projection.
B.3 Energy inequality on the orthogonal part
Take the $L^{2}_{x}L^{2}_{v}(M^{-1})$ inner product of the $R_{\varepsilon}^{\perp}$ equation with $R_{\varepsilon}^{\perp}$. Using coercivity (A5(a)) and integration by parts in $x$ (A2) yields the dissipation inequality
$\frac{1}{2}\frac{d}{dt}\|R_{\varepsilon}^{\perp}\|^{2}+\frac{\lambda_{0}}{\varepsilon}\|R_{\varepsilon}^{\perp}\|^{2}\leq C_{1}\bigl(\|R_{\varepsilon}^{\perp}\|^{2}+\|R_{\varepsilon}^{\parallel}\|^{2}\bigr)+C_{2}\varepsilon^{4},$ (B.1)
where $C_{1}$ depends on macroscopic $H^{s}_{x}$ norms and collision kernel parameters, and $C_{2}$ arises from $\mathcal{S}_{\varepsilon}$ which is $O(\varepsilon^{2})$ in amplitude and contributes $O(\varepsilon^{4})$ to the energy balance.
B.4 Moment equations for the parallel part
Projecting onto $N(L_{x})$ and using conservation laws yields a closed finite-dimensional ODE system for the moments composing $R_{\varepsilon}^{\parallel}$. Schematically,
$\frac{d}{dt}R_{\varepsilon}^{\parallel}=\mathcal{M}R_{\varepsilon}^{\parallel}+\mathcal{F}_{\varepsilon},$
where $\mathcal{M}$ depends on macroscopic gradients and $\mathcal{F}_{\varepsilon}$ is $O(\varepsilon^{2})$ in the $L^{2}_{x}$ norm (coming from quadratic residuals and projections of $\mathcal{S}_{\varepsilon}$). Solving the ODE or using energy bounds gives
$\|R_{\varepsilon}^{\parallel}(t)\|\leq C_{3}\bigl(\|R_{\varepsilon}^{\parallel}(0)\|+\varepsilon^{2}\bigr)e^{C_{4}t},$ (B.2)
with constants $C_{3},C_{4}$ depending on uniform macroscopic norms.
B.5 Grönwall closure
Combine (B.1) and (B.2) to obtain
$\frac{d}{dt}\|R_{\varepsilon}^{\perp}\|^{2}+\frac{2\lambda_{0}}{\varepsilon}\|R_{\varepsilon}^{\perp}\|^{2}\leq C_{5}\bigl(\|R_{\varepsilon}^{\perp}\|^{2}+\varepsilon^{4}+(\|R_{\varepsilon}^{\parallel}(0)\|+\varepsilon^{2})^{2}e^{2C_{4}t}\bigr).$
For $\varepsilon$ sufficiently small the dissipative term $\tfrac{2\lambda_{0}}{\varepsilon}\|R_{\varepsilon}^{\perp}\|^{2}$ dominates the RHS term proportional to $\|R_{\varepsilon}^{\perp}\|^{2}$, allowing absorption and yielding an estimate of the form
$\|R_{\varepsilon}^{\perp}(t)\|^{2}\leq C_{6}\varepsilon^{4}+C_{7}\|R_{\varepsilon}^{\parallel}(0)\|^{2}e^{C_{8}t},$
and using (B.2) with small initial parallel remainder $O(\varepsilon^{2})$ one obtains the uniform-in-time bound
$\|R_{\varepsilon}(t)\|_{L^{2}_{x}L^{2}_{v}(M^{-1})}\leq C\varepsilon^{2},\qquad t\in[0,T],$
with $C$ depending on kernel parameters and macroscopic $H^{s}_{x}$ norms but independent of $\varepsilon$ for $\varepsilon<\varepsilon_{0}$. The hypocoercive alternative follows from replacing the $L^{2}$ energy functional by Villani’s hypocoercive Lyapunov functional (kinetic energy plus mixed derivative corrections); the same Grönwall closure holds with constants tracking derivative losses [15, 14].